Complete the table to the right.
Grade on Business Statistics Exam Frequency RelativeFrequency
?A: 90-100 0.06
?B: 80-89 72
?C: 65-79 192
?D: 50-64 64
?F: Below 50 48
Total 400
Grade on Business Statistics Exam                   Frequency                         Relative Frequency
A: 90-100                                                                          ?                                                   0.06
Answer #1

Frequency fpr Grade ?A: 90-100 = 400-72-192-64-48 = 24